em 7 Constants I Periodic Table Part A A Doppler blood flow unilt emits ultrasound at 5.4 MH What is the frequency shift of the ultrasound reflected from blood moving in an artery at a speed of 023 m/s? Express your answer with the appropriate units Δ/-1 Value Units
Answer #1

As the blood flow unit is supposed to be a receiver at first emission and then has to become a source by reflecting it back,

f'=f*\frac{v-v_{o}}{v+v_{s}}

f'=5.4*\frac{1540-0.23}{1540+0.23}

f'=5.398387 MHz

Shift in frequency

\Delta f = f'-f

\Delta f = 5..398387-5.4

\Delta f = 1.612 kHz

You can also use a direct formula,

\frac{\Delta f}{f} =\frac{2v_{target}}{v}

You should get the same answer